# Studying magnetic circular vortex dichroism effect for photoionization of Rydberg atoms with vortex photons

### v1

@@ II.3 Differential Cross Section
where $E_{f}$ is the kinetic energy of the final electron, $E_{i}$ is the binding energy of the initial electron, $\omega$ is the energy of the incident photon, and $j$ represents the incident photon current. For different initial photon state, current $j$ will be different.
@@ III.1 Magnetic circular-vortex dichroism effect
The magnetic circular dichroism (MCD) effect manifests distinct absorption or scattering probabilities for left- and right-circularly polarized photons incident on chiral targets in a static magnetic field. This phenomenon has emerged as a powerful tool for probing the magnetic properties of materials. For vortex photons, the analogous magnetic vortex dichroism (MVD) effect arises from different absorption or scattering of optical vortices carrying opposite orbital angular momentum (or topological charge). We introduce a novel dichroic phenomenon, termed magnetic circular-vortex dichroism (MCVD), which characterizes the different response of a magnetized target to polarized vortex beams with opposite TAM and helicity:
$\mathrm{MCVD}=\frac{\sigma(m_{\gamma},1)-\sigma(-m_{\gamma},-1)}{\sigma(m_{\gamma},1)+\sigma(-m_{r},-1)}$}.
@@
The underlying symmetry corresponds to a mirror inversion in the transverse plane, equivalent to a one-dimensional parity transformation. Under this operation, factors in Eq. II.2.3:
@@
$\mathrm{MCVD}(m_{\gamma},m_{l})$ $=\frac{\sigma(m_{\gamma},1,{m_{l}})-\sigma(-m_{\gamma},-1,{m_{l}})}{\sigma(m_{\gamma},1,{m_{l}})+\sigma(-m_{\gamma},-1,{m_{l}})}$
$=\frac{\sigma(m_{\gamma},1,{m_{l}})-\sigma(m_{\gamma},1,-{m_{l}})}{\sigma(m_{\gamma},1,{m_{l}})+\sigma(m_{\gamma},1,-{m_{l}})}.$ (20)
@@
III.2 $m$ dependence of MCVD for Rydberg Atoms
To investigate the sensitivity of vortex photons to the magnetic moment of Rydberg atoms, we analyze the dependence of MCVD on the magnetic quantum number $m_{l}$} under two distinct experimental configurations. In both cases, we consider several types of vortex photons characterized by different TAM quantum number $m_{\gamma}$ and helicity $\Lambda_{\gamma}$}.
Figure 1: MCVD in Rydberg atom photoionization with vortex photons. The principal and orbital quantum numbers are fixed at $n=l+1=51$, while the magnetic quantum number varies from $10$ to $50$. Initial photon energy is chosen to be $100$ eV. Conical angle of vortex photon is $\theta_{k}=\pi/10$. Different lines corresponds to different cases with $m_{\gamma}$ ranging from 0 to $\pm 3$.
@@ III.2 $m$ dependence of MCVD for Rydberg Atoms
Figure 2: MCVD in CRA photoionization with vortex photons. Targets follow $n=l+1={m_{l}}+1$ with $n$ varying from $11$ to $51$. Initial photon energy is chosen to be $100$ eV. Conical angle of vortex photon is $\theta_{k}=\pi/10$. Different lines corresponds to different cases with $m_{\gamma}$ ranging from 0 to $\pm 3$.
@@
In the second scenario, we investigate circular Rydberg atoms (CRAs) with varying principal quantum numbers $n$ while maintaining other parameters constant. With the magnetic quantum number fixed as ${m_{l}}=n-1$ (considering only ${m_{l}}\geq 0$), this configuration effectively studies $m_{l}$}-dependence. Figure 2 presents MCVD curves with $m_{l}$} ranging from $10$ to $50$. Again, for MCVD which we have defined in Eq. III.1, we examine vortex photons with TAM quantum number ranging from 0 to $\pm 3$.
@@ III.3 Photon-energy dependence of MCVD for Rydberg Atoms
Figure 3: MCVD in Rydberg atom photoionization with vortex photons. Targets follow $n=l+1=51$ and ${m_{l}}=20$. $E_{k}$ is energy of initial photon ranging from 10 eV to 100 eV. Conical angle of vortex photon is $\theta_{k}=\pi/10$. Different lines corresponds to different cases with $m_{\gamma}$ ranging from 0 to $\pm 3$.
In addition to the magnetic moment of Rydberg atoms, the MCVD is also influenced by the initial photon energy. Figures 3 and 4 demonstrate this dependence for the two configurations defined previously, using vortex photons with TAM quantum numbers $m_{\gamma}=0,\pm 1,\pm 2,\pm 3$.
@@
Figure 4: MCVD in CRA photoionization with vortex photons. Targets follow $n=l+1={m_{l}}+1$ with $n=51$. $E_{k}$ is energy of initial photon ranging from 10 eV to 100 eV. Conical angle of vortex photon is $\theta_{k}=\pi/10$. Different lines corresponds to different cases with $m_{\gamma}$ ranging from 0 to $\pm 3$.

### v2

@@ II.4 Differential Cross Section
[…]
@@
II.5 Multipole expansion and approximations
In the next section, we will demonstrate a range of phenomena using the exact calculations described above. Prior to that, however, it will be useful to introduce an approximation method based on multipole expansion, which will aid in analyzing the upcoming phenomena.
II.5.1 multipole expansion
We can perform a multipole expansion for the photon field (see formula $(16.58)$ in reference [17]):
${{\rm e}^{i\mathbf{k}\cdot\mathbf{r}}=\sum_{{{\eta}}=0}^{\infty}\frac{1}{{{\eta}}!}(i\mathbf{k}\cdot\mathbf{r})^{{{\eta}}}.}$ (21)
Then Eq. (3) becomes
$\mathcal{M}_{fi}^{P}(\mathbf{k},\mathbf{p})$ $=-N_{B}\frac{e}{m_{e}}\sum_{{{\eta}}=0}^{\infty}\Bigg[\frac{1}{{{\eta}}!}\int d^{3}\mathbf{r}\,e^{-i\mathbf{p}\cdot\mathbf{r}}(i\mathbf{k}\cdot\mathbf{r})^{{{\eta}}}$
$=-N_{B}\frac{e}{m_{e}}\sum_{{{\eta}}=0}^{\infty}\left[\frac{1}{{{\eta}}!}(\mathbf{\epsilon}_{k{\Lambda_{\gamma}}}\cdot\mathbf{p})(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{{{\eta}}}\tilde{\psi}_{i}(\mathbf{p})\right],$ (22)
where
[…]
The transition matrix for the vortex case is
$\mathcal{M}^{V}_{fi}(\mathbf{p})$ $=N_{B}\int\frac{d\phi_{\mathbf{k}}}{2\pi}e^{im_{\gamma}\phi_{\mathbf{k}}}\mathcal{M}^{P}_{fi}(\mathbf{k})$
$=-N_{B}\frac{e}{m_{e}}\sum_{{{\eta}}=0}^{\infty}\Bigg[\frac{1}{{{\eta}}!}\int\frac{d\phi}{2\pi}e^{im_{\gamma}\phi_{\mathbf{k}}}(\mathbf{\epsilon}_{k{\Lambda_{\gamma}}}\cdot\mathbf{p})$
$\quad\times(-\mathbf{k}\cdot\mathbf{\nabla}_{p})^{{{\eta}}}\tilde{\psi}_{i}(\mathbf{p})\Bigg]$
$=-N_{B}\frac{e}{m_{e}}e^{im_{\gamma}\phi_{\mathbf{p}}}\sum_{{{\eta}}=0}^{\infty}\Bigg[\frac{1}{{{\eta}}!}\int\frac{d\phi}{2\pi}e^{im_{\gamma}\phi}(\mathbf{\epsilon}_{k{\Lambda_{\gamma}}}\cdot\mathbf{p})$
$\quad\times(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{{{\eta}}}\tilde{\psi}_{i}(\mathbf{p})\Bigg].$ (24)
By considering Eq. (II.2), (II.5.1), and (II.5.1), we can analyze the dependence of multipole contributions on TAM of the vortex photon. For the $\eta$}}-th order multipole contribution:
  • The factor $(-\mathbf{k}\cdot\mathbf{\nabla}_{p})^{{{\eta}}}$ contains $\phi$-dependent phase terms $e^{im_{\nabla}\phi}$ with $m_{\nabla}=0,~\pm 1,~\cdots,~\pm{{\eta}}$
The global $\phi$-dependent phase factor $e^{i(m_{\gamma}+m_{\Lambda}+m_{\nabla})\phi}$ serves as a filter when integrated over one complete cycle. This integration selects only terms satisfying the conservation condition: $m_{\gamma}+m_{\Lambda}+m_{\nabla}=0$, which constrains the photon TAM to $m_{\gamma}=-m_{\Lambda}-m_{\nabla}$, with allowed values spanning 0 to $\pm({{\eta}}+1)$. The multipole expansion exhibits the following characteristic hierarchy:
  • Dipole(${{\eta}}=0$): $m_{\gamma}\in\{0,\pm 1\}$
  • Quadrupole(${{\eta}}=1$): $m_{\gamma}\in\{0,\pm 1,\pm 2\}$
  • Hexapole(${{\eta}}=2$): $m_{\gamma}\in\{0,\pm 1,\pm 2\,\pm 3\}$
This pattern extends to higher multipole orders. For photo-ionization with vortex photon at fixed $m_{\gamma}$, the minimal contributing order is ${{\eta}}=|m_{\gamma}|-1$ for $|m_{\gamma}|\geq 1$, while the $m_{\gamma}=0$ case begins with ${{\eta}}=0$.
II.5.2 Approximations
Since the operator $\partial_{p}$ always yields a $1/p$ term when acting on $\tilde{\psi}_{nlm_{l}}$, we can extract a global factor $k/p$ from Eq. II.5.1. This factor is typically small for non-relativistic photoionization processes. Consequently, higher-order terms generally has smaller contributions to Eq. II.5.1.
With the $\phi$ integral, we explicitly get $m_{\nabla}=-m_{\gamma}-m_{\Lambda}$. The lowest order contribution has ${{\eta}}=|m_{\nabla}|=|m_{\gamma}+m_{\Lambda}|$. Eq. II.5.1 contains $\phi$-independent terms that has no contribution to the lowest order term, due to the filtering effect of $\phi$ integral. We can therefore neglect these terms when evaluating Eq. II.5.1, leading to the simplified expression:
[…]
For Rydberg atoms, the wave function in momentum space is given by Eq. II.3.1:
$\tilde{\psi}_{nl{m_{l}}}(\mathbf{p})\propto\frac{Y_{l}^{m_{l}}(\theta_{\mathbf{p}},\phi_{\mathbf{p}})}{n^{2(l+2)}a^{2(l+2)}p^{l+4}}.$ (27)
[…]
$-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}}$ $\approx\frac{k\sin\theta_{\mathbf{k}}}{p}\Big\{\Big[(l+2)\sin\theta_{\mathbf{p}}-\frac{l}{\sin\theta_{\mathbf{p}}}\Big]e^{i\lambda\phi}$
$\quad+(l+2)\sin\theta_{\mathbf{p}}e^{-i\lambda\phi}\Big\},$
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{2}$ $\approx\frac{k^{2}\sin^{2}\theta_{\mathbf{k}}}{p^{2}}\Big\{\Big[(l+2)\sin\theta_{\mathbf{p}}-\frac{l}{\sin\theta_{\mathbf{p}}}\Big]e^{i\lambda\phi}$
$\quad+(l+2)\sin\theta_{\mathbf{p}}e^{-i\lambda\phi}\Big\}\Big\{\Big[(l+\frac{5}{2})\sin\theta_{\mathbf{p}}$
$\quad-\frac{l}{\sin\theta_{\mathbf{p}}}\Big]e^{i\lambda\phi}+(l+\frac{5}{2})\sin\theta_{\mathbf{p}}e^{-i\lambda\phi}\Big\}$
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{{{\eta}}}$ $\approx\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{{{\eta}}}\prod_{j=1}^{{{\eta}}}\Big\{\Big[(l+\frac{3+j}{2})\sin\theta_{\mathbf{p}}$
$\quad-\frac{l}{\sin\theta_{\mathbf{p}}}\Big]e^{i\lambda\phi}+(l+\frac{3+j}{2})\sin\theta_{\mathbf{p}}e^{-i\lambda\phi}\Big\}.$ (28)
Specially for CRA, the electron wavefunction concentrates near $\theta_{\mathbf{p}}\approx\pi/2$, we obtain further approximations for the multipole operators:
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{{{\eta}}}$ $\approx\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{{{\eta}}}\prod_{j=1}^{{{\eta}}}\Bigg[\left(l+\frac{j+3}{2}\right)e^{-i\lambda\phi}$
$\quad+\left(\frac{j+3}{2}\right)e^{i\lambda\phi}\Bigg].$ (29)
The higher-order contributions are strongly suppressed by the $(k\sin\theta_{\mathbf{k}}/p)^{{\eta}}$ factor, with the $\sin\theta_{\mathbf{k}}$ term providing additional suppression.
@@ III.1 Magnetic circular-vortex dichroism effect
The magnetic circular dichroism (MCD) effect manifests distinct absorption or scattering probabilities for left- and right-circularly polarized photons incident on chiral targets in a static magnetic field. This phenomenon has emerged as a powerful tool for probing the magnetic properties of materials. For vortex photons, the analogous magnetic vortex dichroism (MVD) effect arises from different absorption or scattering of optical vortices carrying opposite OAM (or topological charge). We introduce a novel dichroic phenomenon, termed magnetic circular-vortex dichroism (MCVD), which characterizes the different response of a magnetized target to polarized vortex beams with opposite TAM and helicity:
${\mathrm{MCVD}=\frac{\sigma(m_{\gamma},\Lambda_{\gamma}=1)-\sigma(-m_{\gamma},\Lambda_{\gamma}=-1)}{\sigma(m_{\gamma},\Lambda_{\gamma}=1)+\sigma(-m_{r},\Lambda_{\gamma}=-1)}}.$ (30)
It is a physical observable measured via a two-collision process, in which vortex beams with opposite TAM and helicity are directed at the same polarized atom. And it is the core subject of this work. For simplicity, in the following text, we use $(m_{\gamma},\Lambda_{\gamma})$” to refer to the MCVD in different initial parameters defined in Eq. 30.
Note that we use the eigenstates quantized by four quantum numbers: energy, longitudinal momentum, TAM and helicity. OAM is not a good quantum number in general. Only for the paraxial case that OAM can approximately be a good quantum number. The two states involving in MCVD has opposite TAM and helicity. Their OAM is determined by relation ”TAM=OAM+helicity” for paraxial case.
@@
The underlying symmetry corresponds to a mirror inversion in the transverse plane, equivalent to a one-dimensional parity transformation. Under this operation, factors in Eq. II.3.3:
@@
$\mathrm{MCVD}$ $=\frac{\sigma(m_{\gamma},\Lambda_{\gamma}=1,{m_{l}})-\sigma(-m_{\gamma},\Lambda_{\gamma}=-1,{m_{l}})}{\sigma(m_{\gamma},\Lambda_{\gamma}=1,{m_{l}})+\sigma(-m_{\gamma},\Lambda_{\gamma}=-1,{m_{l}})}$
$=\frac{\sigma(m_{\gamma},\Lambda_{\gamma}=1,{m_{l}})-\sigma(m_{\gamma},\Lambda_{\gamma}=1,-{m_{l}})}{\sigma(m_{\gamma},\Lambda_{\gamma}=1,{m_{l}})+\sigma(m_{\gamma},\Lambda_{\gamma}=1,-{m_{l}})}.$ (31)
@@
III.2 $m_{l}$ dependence of MCVD for Rydberg Atoms
To investigate the sensitivity of vortex photons to the magnetic moment of Rydberg atoms, we analyze the dependence of MCVD on the magnetic quantum number $m_{l}$} under two distinct experimental configurations. In both cases, we consider several types of vortex photons characterized by different TAM quantum number $m_{\gamma}$ and helicity $\Lambda_{\gamma}$}. A Rydberg atom with main quantum number $n=51$ is investigated. Its radius is about 138138nm. Initial vortex photon has energy $\omega=100$eV and conical angle $\theta_{k}=\pi/10$. In these parameters, the transverse characteristic length for vortex photon is about 365365nm which is in the same order with Rydberg atom’s radius.
Figure 4: MCVD in Rydberg atom photoionization with vortex photons. The principal and orbital quantum numbers are fixed at $n=l+1=51$, while the magnetic quantum number varies from $10$ to $50$. Initial photon energy is chosen to be $100$ eV. Conical angle of vortex photon is $\theta_{k}=\pi/10$. Different lines corresponds to different cases with $m_{\gamma}$ ranging from 0 to $\pm 3$. Label $(a,b)$ in the legends means we get MCVD from difference of two initial cases: $(m_{\gamma}=a,\Lambda_{\gamma}=b)$ and ($m_{\gamma}=-a,\Lambda_{\gamma}=-b)$.
@@ III.2 $m_{l}$ dependence of MCVD for Rydberg Atoms
Figure 5: MCVD in CRA photoionization with vortex photons. Targets follow $n=l+1={m_{l}}+1$ with $n$ varying from $11$ to $51$. Initial photon energy is chosen to be $100$ eV. Conical angle of vortex photon is $\theta_{k}=\pi/10$. Different lines corresponds to different cases with $m_{\gamma}$ ranging from 0 to $\pm 3$. Label $(a,b)$ in the legends means we get MCVD from difference of two initial cases: $(m_{\gamma}=a,\Lambda_{\gamma}=b)$ and ($m_{\gamma}=-a,\Lambda_{\gamma}=-b)$.
@@
In the first configuration, the shape of the wave function varies with the magnetic quantum number $m_{l}$. In contrast, the second scenario employs CRAs with different principal quantum numbers $n$ to isolate the effect of $m_{l}$ under a unchanging wave function shape. With the magnetic quantum number fixed as ${m_{l}}=n-1$ (considering only ${m_{l}}\geq 0$), this configuration effectively studies $m_{l}$}-dependence. Figure 5 presents MCVD curves with $m_{l}$} ranging from $10$ to $50$. Again, for MCVD which we have defined in Eq. III.1, we examine vortex photons with TAM quantum number ranging from 0 to $\pm 3$.
@@
phenomenon analyzing
As established earlier, the $\phi$ integral in Eq. II.5.1 acts as a selection filter through the condition $m_{\gamma}+m_{\Lambda}+m_{\nabla}=0$, where $m_{\Lambda}=0$ or $\pm 1$ originates from the polarization term:
The phase contributions $m_{\nabla}$ emerge from the operator $(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{{{\eta}}}$.
When $|m_{\gamma}|\geq 2$, the dipole contribution (${{\eta}}=0$) vanishes, and the dominant contribution arises from ${{\eta}}=|m_{\nabla}|\geq 1$. According to Eq. II.5.2, four distinct cases emerge:
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{|m_{\nabla}|}$ $\rightarrow\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{|m_{\nabla}|}e^{-i|m_{\nabla}|\phi}$
$\quad\times\prod_{j=0}^{|m_{\nabla}|-1}\Big[\left(l+2+\frac{j}{2}\right)\sin\theta_{\mathbf{p}}\Big]$ (33)
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{|m_{\nabla}|}$ $\rightarrow\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{|m_{\nabla}|}e^{-i|m_{\nabla}|\phi}$
$\quad\times\prod_{j=0}^{|m_{\nabla}|-1}\Big[\left(l+2+\frac{j}{2}\right)\sin\theta_{\mathbf{p}}-\frac{l}{\sin\theta_{\mathbf{p}}}\Big]$ (34)
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{|m_{\nabla}|}$ $\rightarrow\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{|m_{\nabla}|}e^{i|m_{\nabla}|\phi}$
$\quad\times\prod_{j=0}^{|m_{\nabla}|-1}\Big[\left(l+2+\frac{j}{2}\right)\sin\theta_{\mathbf{p}}\Big]$ (35)
$(-\mathbf{k}\cdot\mathbf{\nabla}_{\mathbf{p}})^{|m_{\nabla}|}$ $\rightarrow\left(\frac{k\sin\theta_{\mathbf{k}}}{p}\right)^{|m_{\nabla}|}e^{i|m_{\nabla}|\phi}$
$\quad\times\prod_{j=0}^{|m_{\nabla}|-1}\Big[\left(l+2+\frac{j}{2}\right)\sin\theta_{\mathbf{p}}-\frac{l}{\sin\theta_{\mathbf{p}}}\Big]$ (36)
When $m_{l}$ is close to $l$, the electronic wave function is predominantly localized near $\theta_{\mathbf{p}}=\pi/2$. In this region, the term $(l+2+j/2)\sin\theta_{\mathbf{p}}$ significantly exceeds $(l+2+j/2)\sin\theta_{\mathbf{p}}-l/\sin\theta_{\mathbf{p}}$. As a result, when $|m_{\gamma}|$ and $|m_{l}|$ are fixed, photoionization favors vortex photons that co-rotate with the atomic angular momentum (i.e., $m_{\gamma}>0$ for $m_{l}>0$, and $m_{\gamma}<0$ for $m_{l}<0$). This selectivity leads to MCVD values approaching $\pm 1$, as shown for large $m_{l}$ in Fig. 4 and for all $m_{l}$ in Figs. 5 when $|m_{\gamma}|=2$ and $|m_{\gamma}|=3$. These observations reveal the chiral nature of vortex-photon photoionization from polarized Rydberg atoms. Conversely, when $m_{l}$ is much smaller than $l$, the electronic wave function is distributed across the entire angular range $(0,\pi)$. Here, $(l+2+j/2)\sin\theta_{\mathbf{p}}$ is no longer dominant over $(l+2+j/2)\sin\theta_{\mathbf{p}}-l/\sin\theta_{\mathbf{p}}$, and processes with both co-rotating and counter-rotating vortex photons ($m_{\gamma}$ and $m_{l}$ having the same or opposite signs) become comparably important. Consequently, the MCVD values lie between 0 and $\pm 1$, as seen for small $m_{l}$ with $|m_{\gamma}|=2$ and $|m_{\gamma}|=3$ in Fig. 4.
When $|m_{\gamma}|=1$, the dipole contribution (${{\eta}}=0$) plays important role. Two scenarios occur:
1. When $m_{\gamma}$ and $\Lambda_{\gamma}$} share the same sign, the $p_{\perp}\cos^{2}(\theta_{\mathbf{k}}/2)e^{-i{\Lambda_{\gamma}}\phi}$ term dominates, i.e. dipole contribution dominates, yielding MCVD $\approx 0$ resembling initial plane wave case (see $m_{\gamma}=1$ case in Figs. 4 and 5).
2. When $m_{\gamma}$ and $\Lambda_{\gamma}$} have opposite signs, the $p_{\perp}\sin^{2}(\theta_{\mathbf{k}}/2)e^{i{\Lambda_{\gamma}}\phi}$ term, which gives dipole contribution, is suppressed by $\sin^{2}(\theta_{\mathbf{k}}/2)$. ${{\eta}}=1,2$ contributions (given by $p_{z}\sin\theta_{k}$ and $p_{\perp}\cos^{2}(\theta_{\mathbf{k}}/2)e^{-i{\Lambda_{\gamma}}\phi}$ terms respectively) become important. Their interference with dipole contribution introduce significant $m_{l}$}-dependent MCVD (see $m_{\gamma}=-1$ case in Figs. 4 and 5).
When $m_{\gamma}=0$, the dipole contribution is given by $p_{z}\sin\theta_{k}$ term and suppressed by $\sin\theta_{k}$. The quadrupole contribution (${{\eta}}=1$) given by $p_{\perp}\cos^{2}(\theta_{\mathbf{k}}/2)e^{-i{\Lambda_{\gamma}}\phi}$ term can be comparable to the dipole contribution. Their interference introduce significant $m_{l}$}-dependent MCVD (see $m_{\gamma}=0$ case in Figs. 4 and 5).
@@ III.3 Photon-energy dependence of MCVD for Rydberg Atoms
Figure 6: MCVD in Rydberg atom photoionization with vortex photons. Targets follow $n=l+1=51$ and ${m_{l}}=20$. $E_{k}$ is energy of initial photon ranging from 10 eV to 100 eV. Conical angle of vortex photon is $\theta_{k}=\pi/10$. Different lines corresponds to different cases with $m_{\gamma}$ ranging from 0 to $\pm 3$. Label $(a,b)$ in the legends means we get MCVD from difference of two initial cases: $(m_{\gamma}=a,\Lambda_{\gamma}=b)$ and ($m_{\gamma}=-a,\Lambda_{\gamma}=-b)$.
In addition to the magnetic moment of Rydberg atoms, the MCVD is also influenced by the initial photon energy. Figures 6 and 7 demonstrate this dependence for the two configurations defined previously, using vortex photons with TAM quantum numbers $m_{\gamma}=0,\pm 1,\pm 2,\pm 3$.
@@
Figure 7: MCVD in CRA photoionization with vortex photons. Targets follow $n=l+1={m_{l}}+1$ with $n=51$. $E_{k}$ is energy of initial photon ranging from 10 eV to 100 eV. Conical angle of vortex photon is $\theta_{k}=\pi/10$. Different lines corresponds to different initial energy cases with $m_{\gamma}$ ranging from 0 to $\pm 3$. Label $(a,b)$ in the legends means we get MCVD from difference of two initial cases: $(m_{\gamma}=a,\Lambda_{\gamma}=b)$ and ($m_{\gamma}=-a,\Lambda_{\gamma}=-b)$.
@@
phenomenon analyzing
These phenomena can be readily understood. As shown earlier, when $|m_{\gamma}|\geq 2$, the dominant multipole contribution has order ${{\eta}}=|m_{\gamma}|-1$ and scales as $k^{|m_{\gamma}|-1}$. As a result, the MCVD becomes independent of $k$. Since $E_{k}=k$ for massless photons, it follows that the MCVD is also independent of the photon energy $E_{k}$, as illustrated in Fig. 6 and Fig. 7 for $|m_{\gamma}|\geq 2$. For $m_{\gamma}=0$ or $m_{\gamma}=-1$, both dipole and quadrupole contributions are significant. The dipole term is independent of $k$, while the quadrupole term scales linearly with $k$. This leads to an $E_{k}$-dependent MCVD, as seen in Fig. 6 and Fig. 7 for these cases. When $m_{\gamma}=1$, the dominant contribution comes from the dipole term, which is independent of $k$. Consequently, the MCVD remains close to zero, as shown in Fig. 6 and Fig. 7 for $m_{\gamma}=1$.